# AnyAttack: Towards Large-scale Self-supervised Generation of Targeted Adversarial Examples for Vision-Language Models

### v1

@@ Abstract
Due to their multimodal capabilities, Vision-Language Models (VLMs) have found numerous impactful applications in real-world scenarios. However, recent studies have revealed that VLMs are vulnerable to image-based adversarial attacks, particularly targeted adversarial images that manipulate the model to generate harmful content specified by the adversary. Current attack methods rely on predefined target labels to create targeted adversarial attacks, which limits their scalability and applicability for large-scale robustness evaluations. In this paper, we propose AnyAttack, a self-supervised framework that generates targeted adversarial images for VLMs without label supervision, allowing any image to serve as a target for the attack. To address the limitation of existing methods that require label supervision, we introduce a contrastive loss that trains a generator on a large-scale unlabeled image dataset, LAION-400M dataset, for generating targeted adversarial noise. This large-scale pre-training endows our method with powerful transferability across a wide range of VLMs. Extensive experiments on five mainstream open-source VLMs (CLIP, BLIP, BLIP2, InstructBLIP, and MiniGPT-4) across three multimodal tasks (image-text retrieval, multimodal classification, and image captioning) demonstrate the effectiveness of our attack. Additionally, we successfully transfer AnyAttack to multiple commercial VLMs, including Google’s Gemini, Claude’s Sonnet, and Microsoft’s Copilot. These results reveal an unprecedented risk to VLMs, highlighting the need for effective countermeasures.
@@ 1 Introduction
Vision-Language Models (VLMs) have exhibited remarkable performance across a diverse array of tasks, primarily attributed to the scale of training data and model size (Radford et al., 2021; Li et al., 2023; Zhu et al., 2024). Despite their remarkable performance, these models, heavily reliant on visual inputs, remain vulnerable to image-based adversarial attacks [^1] For simplicity, we will refer to image-based adversarial attacks as “adversarial attacks” in the remainder of this paper, distinguishing them from text-based adversarial attacks., which are carefully crafted input images designed to mislead the model into making incorrect predictions (Szegedy et al., 2013). While general, untargeted adversarial attacks only aim to induce incorrect outputs, targeted adversarial attacks present a more insidious threat, manipulating the model’s output to yield an adversary-specified, predetermined response. For instance, a benign image such as a landscape could be subtly altered to elicit harmful text descriptions such as “violence” or “explicit content” from the model. Such manipulation could have severe implications for content moderation systems, potentially leading to the removal of legitimate content or the inappropriate distribution of harmful material.
@@
Both targeted adversarial attacks and jailbreak techniques (Zou et al., 2023; Bagdasaryan et al., 2023; Carlini et al., 2024) aim to induce harmful responses from models. However, the key distinction lies in their outputs: targeted adversarial attacks produce adversary-specified, predetermined responses, whereas jailbreak attacks elicit non-predetermined responses. We will further elaborate on this distinction in Section 2. As VLMs become increasingly accessible to the public, facilitating the rapid proliferation of downstream applications, this vulnerability poses a significant threat to the reliability and security of VLMs in real-world scenarios. Therefore, exploring new targeted attack methods tailored to VLMs is crucial to address these vulnerabilities. However, existing targeted attack methods on VLMs present challenges due to the reliance on target labels for supervision, which limits the scalability of the training process. For example, it is impractical to expect a generator trained on ImageNet (Russakovsky et al., 2015) to produce effective adversarial noise for VLMs.
To address this limitation, we propose a novel self-supervised learning framework called AnyAttack, which utilizes the original image itself as supervision, allowing any image to serve as a target for the deployment of targeted adversarial attacks. Figure 1 illustrates the differences between our framework and the existing methods. Drawing inspiration from popular contrastive learning techniques (Chen et al., 2020; He et al., 2020; Radford et al., 2021; Li et al., 2023), we introduce a contrastive loss to allow training a generator on the large-scale dataset to generate targeted adversarial noise. This encourages the generated noise to emulate the original image itself (positive pairs) in the embedding space while differentiating from other images (negative pairs). The unsupervised nature of our method enables large-scale pre-training of our generator on the LAION-400M dataset (Schuhmann et al., 2021), allowing for the generation of effective adversarial noise. Benefiting from this large-scale pre-training and exposure to a broader spectrum of images, our method shows powerful transferability across a wide range of VLMs. To the best of our knowledge, this is the first time such a large-scale dataset has been employed to train a generator for the generation of targeted adversarial noise. Notably, the generator can produce adversarial images with any given image, imposing no restrictions on the image itself, thereby introducing an unprecedented security risk to the entire community.
@@
Furthermore, to enable adaptation to specific domains and multimodal tasks, we fine-tune the pre-trained generator on downstream datasets for adapting downstream vision-language tasks. To demonstrate the effectiveness of our approach, we conduct extensive experiments on 5 target VLMs (CLIP (Radford et al., 2021), BLIP (Li et al., 2022), BLIP2 (Li et al., 2023), InstructBLIP (Dai et al., 2023), and MiniGPT-4 (Zhu et al., 2024)), across 3 multimodal tasks (image-text retrieval, multimodal classification, and image captioning). We also evaluate our method on commercial VLMs, including Google’s Gemini, Claude’s Sonnet, and Microsoft’s Copilot.
@@
  • We propose AnyAttack, a novel self-supervised framework for generating targeted adversarial attacks on VLMs without the need for target labels, allowing any image to be used as a target for the attack.
  • We introduce a contrastive loss that facilitates the training of the generator for adversarial noise using unlabeled, large-scale datasets. This makes our AnyAttack framework scalable, effectively overcoming the limitations of previous methods.
  • We demonstrate the effectiveness of our AnyAttack on five mainstream open-source VLMs across three multimodal tasks. Additionally, we successfully transfer our attack to three commercial VLMs. These results offer valuable insights into the vulnerabilities of state-of-the-art models in real-world applications.
Figure 1: (a) Existing methods generate targeted adversarial attacks guided by the target labels. (b) Our AnyAttack is a self-supervised attack method that does not require any labels. The dashed lines highlight the key differences between the two approaches.
@@ 2 Related Work
Targeted Adversarial Attacks
@@
A number of works have been proposed to enhance the effectiveness and transferability of targeted adversarial attacks against vision models. Input augmentation techniques like image translation (Dong et al., 2019), cropping (Wei et al., 2023), mixup (Wang et al., 2021; Liu & Lyu, 2024), and resizing (Xie et al., 2019), have been employed to increase the diversity of adversarial inputs, thereby improving their transferability across different target models. Additionally, adversarial fine-tuning and model enhancement techniques have been explored to bolster the attack capabilities of surrogate models (Springer et al., 2021; Zhang et al., 2023; Wu et al., 2024). These methods typically involve retraining the surrogate models with a mix of clean and adversarial examples to improve their robustness against future attacks. Furthermore, optimization techniques have evolved to stabilize the update processes during adversarial training. Methods such as adaptive learning rates and gradient clipping have been integrated to ensure more consistent updates and enhance the overall performance of the adversarial attacks (Dong et al., 2018; Wang & He, 2021; Lin et al., 2023). These advancements collectively contribute to the development of more effective and transferable adversarial attacks in the realm of vision models.
Jailbreak Attacks on VLMs
@@
VLMs have revolutionized DNNs by leveraging large-scale pre-training on diverse image-text datasets. These models learn to integrate visual and textual information effectively, enabling superb performance across a wide range of tasks. Broadly, VLMs can be categorized into two types: the first offers multimodal functionalities built on large language models (LLMs), complemented by visual models, such as BLIP2 (Li et al., 2023), InstructBLIP (Dai et al., 2023), and MiniGPT-4 (Zhu et al., 2024). The second type provides a more balanced approach, bridging textual and visual modalities efficiently, as seen in models like CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), and BLIP (Li et al., 2022). Recent advancements in VLMs have spurred parallel research into their vulnerabilities, with jailbreaks and adversarial attacks emerging as distinct threat vectors. Multimodal jailbreaks primarily exploit cross-modal interaction vulnerabilities in VLMs, with the intention of influencing LLMs (Zou et al., 2023). These attacks manipulate inputs of text (Wu et al., 2023), images (Carlini et al., 2024; Gong et al., 2023; Qi et al., 2024; Niu et al., 2024), or both simultaneously (Ying et al., 2024; Wang et al., 2024), aiming to elicit harmful but non-predefined responses. In contrast, image-based adversarial attacks focus on manipulating the image encoder of VLMs, typically leaving the LLM component largely undisturbed. The objective is to induce adversary-specified, predetermined responses through precise visual manipulations. Understanding these differences is crucial for explaining our methodology.
Adversarial Attacks on VLMs
@@
Adversarial research on VLMs is relatively limited compared to the extensive studies on vision models, with the majority of existing attacks focusing primarily on untargeted attacks. Co-Attack (Zhang et al., 2022) was among the first to perform white-box untargeted attacks on several VLMs. Following this, more approaches have been proposed to enhance adversarial transferability for black-box untargeted attacks (Lu et al., 2023; Zhou et al., 2023; Yin et al., 2024; Xu et al., 2024). Cross-Prompt Attack (Luo et al., 2024) investigates a novel setup for adversarial transferability based on the prompts of language models. AttackVLM (Zhao et al., 2024) is the most closely related work, using a combination of text inputs and popular text-to-image models to generate guided images for creating targeted adversarial images. Although their approach shares a similar objective with our work, our method distinguishes itself by being self-supervised and independent of any text-based guidance.
@@ 3 Proposed Attack
In this section, we first present the preliminaries on targeted adversarial attacks and then introduce our proposed AnyAttack and its two phases (i.e., pre-training and fine-tuning).
3.1 Problem Formulation
Threat Model
@@ 3.1 Problem Formulation
We begin by formulating the problem of targeted adversarial attacks. Let $f_{s}$ represent a pre-trained surrogate model, and $\mathcal{D}=\{(x,y)\}$ denote the image dataset, where $x$ is the original image and $y$ is the corresponding label (description). The attacker’s objective is to craft an adversarial example $x^{\prime}=x+\delta$ that misleads the target model $f_{t}$ into predicting a predefined target label $y_{t}$. In the context of VLMs, this objective requires that $x^{\prime}$ aligns with $y_{t}$ as a valid image-text pair. The process of generating targeted adversarial images typically involves finding a perturbation $\delta$ using the surrogate model $f_{s}$. Existing strategies can be approached through two primary strategies: target label supervision and target image supervision.
Table 1: The supervision and formulation of different targeted adversarial attack strategies. The first two rely on explicit target supervision (label or image), whereas our AnyAttack is unsupervised.
Strategy Formulation Target label supervision $\min\mathcal{L}(f_{s}(x+\delta),f_{s}(y_{t})),\quad\text{s.t.}\ y_{t}\neq y$ Target image supervision $\min\mathcal{L}(f_{s}(x+\delta),f_{s}(x_{t})),\quad\text{s.t.}\ y_{t}\neq y$ AnyAttack (ours) $\min\mathcal{L}(f_{s}(\delta+x_{r}),f_{s}(x)),\quad\text{s.t.}\ x_{r}\neq x$
The first approach utilizes the target label $y_{t}$ as supervision, directing the embedding of the adversarial image $x^{\prime}$ to align with that of $y_{t}$, as demonstrated in AttackVLM-it Zhao et al. (2024). The second approach employs the target image $x_{t}$, which corresponds to $y_{t}$, as supervision to encourage the embedding of $x^{\prime}$ to replicate that of $x_{t}$. This is illustrated in AttackVLM-ii (Zhao et al., 2024) and certain image-based attacks (Wei et al., 2023; Wu et al., 2024). Both methods depend on explicit target supervision, as summarized in Table 1. In these approaches, $\mathcal{L}$ denotes a distance-based loss function, such as Euclidean distance or cosine similarity. In contrast, our method, AnyAttack, employs the input image itself to guide the attack and thus is unsupervised. In this context, $x_{r}$ is a random image that is unrelated to $x$, while the adversarial noise $\delta$ is designed to align with the original image $x$ within the surrogate model’s embedding space. In summary, existing methods generate adversarial noise that mimics other images, whereas our approach produces adversarial noise that closely resembles the original image itself.
@@ 3.2 AnyAttack
Our proposed framework, AnyAttack, employs a two-stage training paradigm: pre-training and fine-tuning. Figure 2 provides a framework overview of AnyAttack.
@@
In the pre-training stage, we train a decoder $F$, to produce adversarial noise $\delta$ on large-scale datasets $\mathcal{D}_{p}$. Given a batch of images $x$, we extract their embeddings using a frozen image encoder $E$. These normalized embeddings $\mathbf{z}$ are then fed into the decoder $F$, which generates adversarial noise $\delta$ corresponding to the images $x$. To enhance generalization and computational efficiency, we introduce a $K$-augmentation strategy that creates multiple shuffled versions of the original images within each mini-batch. During this process, adversarial noise is added to the shuffled original images (unrelated images) to produce the adversarial images. After passing through $E$, we employ a contrastive loss to maximize the cosine similarity between positive sample pairs (the $i$-th elements of the adversarial and original embeddings) while minimizing the similarity between negative sample pairs (the remaining elements). This approach trains the decoder $F$ to ensure that the perturbed images resemble the original images in the embedding space of encoder $E$, while distinguishing them from the shuffled versions.
In the fine-tuning stage, we adapt the pre-trained decoder $F$ to a specific downstream dataset $\mathcal{D}_{f}$. The frozen encoder $E$ continues to provide embeddings that guide the generation of adversarial noise $\delta$. We use an unrelated random image $x_{r}$ from an external dataset $\mathcal{D}_{e}$ as the clean image to synthesize the adversarial image $x_{r}+\delta$. Unlike the pre-training stage, where only the encoder is utilized as a surrogate model, we introduce auxiliary models during the fine-tuning stage to enhance transferability. We tailor various fine-tuning strategies based on the requirements of each task.
Pre-training Stage
Figure 2: An overview of the proposed AnyAttack.
@@
The pre-training phase of AnyAttack aims to train the generator on large-scale datasets, enabling it to handle a diverse array of input images as potential targets. Given a batch of $n$ images $x\in\mathbb{R}^{n\times H\times W\times 3}$ from the large-scale training dataset $\mathcal{D}_{p}$, we employ the CLIP ViT-B/32 image encoder, which is frozen during training, as the encoder $E$, to obtain the normalized embeddings $E(x)=\mathbf{z}\in\mathbb{R}^{n\times d}$ corresponding to the original images $x$, where $d$ represents the embedding dimension (i.e., 512 for CLIP ViT-B/32). Subsequently, we deploy an initialized decoder $F$, which maps the embeddings $\mathbf{z}$ to adversarial noise $D(\mathbf{z})=\delta\in\mathbb{R}^{n\times H\times W\times 3}$ corresponding to the original images $x$. We expect the generated noises $\delta$ to serve as adversarial noise representative of the original images $x$. Our goal is for the generated noises $\delta$, when added to random images $x_{r}$, to be interpreted by the encoder $E$ as the original images $x$, i.e., $E(x_{r}+\delta)=E(x)$.
@@
However, when the number of random images is smaller than the training dataset, the generated noises $\delta$ may overfit to this limited set, leading to poor generalization for $F$. To address this, we propose the $K$-augmentation strategy, which expands the set of random images to match the size of the training dataset $\mathcal{D}_{p}$. This strategy increases the number of sample pairs within each batch by a factor of $K$, thereby improving computational efficiency. Specially, $K$-augmentation duplicates both adversarial noises $\delta$ and the original images $x$ $K$ times, forming $K$ mini-batches. For each mini-batch, the order of the adversarial noises remain consistent, while the order of the original images is shuffled within the mini-batch, referred to as shuffled images. These shuffled images are then added to the corresponding adversarial noise, resulting in adversarial images $x^{\prime}$. Next, the adversarial images are fed into $F$ to produce adversarial embeddings $\mathbf{z}^{(adv)}$, which are then used for subsequent calculations against the original embeddings $\mathbf{z}$. We introduce a contrastive loss that maximizes the cosine similarity between positive sample pairs, defined by the $i$-th elements of adversarial and original embeddings in each mini-batch, while minimizing the similarity between the negative pairs, which consist of all other elements. This setup creates $n$ positive pairs and $n(n-1)$ negative pairs in every mini-batch, with gradients accumulated to update $F$:
$\mathcal{L}_{\text{Con}}=-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\exp\left(\mathbf{z}_{i}\cdot\mathbf{z}_{i}^{(adv)}/\tau(t)\right)}{\sum_{j=1}^{n}\exp\left(\mathbf{z}_{i}\cdot\mathbf{z}_{j}^{(adv)}/\tau(t)\right)},$ (1)
@@
where $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{(adv)}$ are the $\ell_{2}$-normalized embeddings of the $i$-th sample from original images $x$ and adversarial images $x^{\prime}$. $\tau(t)$ is the temperature at step $t$, enabling the model to dynamically adjust the hardness of negative samples during training. To facilitate learning and convergence in early training, we set a relatively large initial temperature $\tau_{0}$ at the beginning of training and gradually decrease it, reaching the final temperature $\tau_{\text{final}}$ after a certain number of steps $T$:
@@
3.2.1 Fine-tuning Stage
In the fine-tuning stage, we refine the pre-trained decoder $F$ on downstream vision-language datasets using task-specific objective functions, facilitating its adaptation to particular domains and multimodal tasks. The motivation for fine-tuning arises from scenarios where well-defined multimodal tasks and in-domain images are available. Given a batch of $n$ images $x\in\mathbb{R}^{n\times H\times W\times 3}$ from the downstream dataset $\mathcal{D}_{f}$, the encoder $E$ remains frozen and outputs the embeddings $\mathbf{z}$, which are then fed into the decoder $F$ to generate the noise $\delta$. We randomly select images from an external dataset $\mathcal{D}_{e}$ as unrelated images $x_{r}$, which are then added to the generated noise $\delta$ to create adversarial images. To improve transferability, we incorporate auxiliary models alongside the encoder $E$, forming an ensemble surrogate. Drawing on research in ensemble learning for adversarial attacks (Liu et al., 2017), we select auxiliary models based on model diversity, as greater differences between models are known to improve complementarity. This ensures that both adversarial and original images maintain consistency across the embedding spaces of multiple models.
Depending on the downstream tasks, we employ two different fine-tuning objectives. The first strategy is tailored for the image-text retrieval task, which imposes stricter requirements for distinguishing between similar samples. It demands robust retrieval performance in both directions: from $\mathbf{z}^{(adv)}$ to $\mathbf{z}$ and from $\mathbf{z}$ to $\mathbf{z}^{(adv)}$. This motivates the adoption of a bidirectional InfoNCE loss:
$\mathcal{L}_{\text{Bi}}=\frac{1}{2n}\sum_{i=1}^{n}\left(-\log\frac{\exp(\mathbf{z}_{i}\cdot\mathbf{z}_{i}^{(adv)}/\tau)}{\sum_{j=1}^{n}\exp(\mathbf{z}_{i}\cdot\mathbf{z}_{j}^{(adv)}/\tau)}-\log\frac{\exp(\mathbf{z}_{i}^{(adv)}\cdot\mathbf{z}_{i}/\tau)}{\sum_{j=1}^{n}\exp(\mathbf{z}_{i}^{(adv)}\cdot\mathbf{z}_{j}/\tau)}\right).$ (3)

### v2

# AnyAttack: Targeted Adversarial Attacks on Vision-Language Models Toward Any Images

@@ Abstract
Due to their multimodal capabilities, Vision-Language Models (VLMs) have found numerous impactful applications in real-world scenarios. However, recent studies have revealed that VLMs are vulnerable to image-based adversarial attacks, particularly targeted adversarial images that manipulate the model to generate harmful content specified by the adversary. Current attack methods rely on predefined target labels to create targeted adversarial attacks, which limits their scalability and applicability for large-scale robustness evaluations. In this paper, we propose AnyAttack, a self-supervised framework that generates targeted adversarial images for VLMs without label supervision, allowing any image to serve as a target for the attack. Our framework employs the “pre-training and fine-tuning” paradigm, with the adversarial noise generator pre-trained on the large-scale LAION-400M dataset. This large-scale pre-training endows our method with powerful transferability across a wide range of VLMs. Extensive experiments on five mainstream open-source VLMs (CLIP, BLIP, BLIP2, InstructBLIP, and MiniGPT-4) across three multimodal tasks (image-text retrieval, multimodal classification, and image captioning) demonstrate the effectiveness of our attack. Additionally, we successfully transfer AnyAttack to multiple commercial VLMs, including Google Gemini, Claude Sonnet, Microsoft Copilot and OpenAI GPT. These results reveal an unprecedented risk to VLMs, highlighting the need for effective countermeasures.
@@ 1 Introduction
Vision-Language Models (VLMs) have exhibited remarkable performance across a diverse array of tasks, primarily attributed to the scale of training data and model size [18, 11, 37]. Despite their remarkable performance, these models, heavily reliant on visual inputs, remain vulnerable to image-based adversarial attacks [^1] For simplicity, we will refer to image-based adversarial attacks as “adversarial attacks” in the remainder of this paper, distinguishing them from text-based adversarial attacks., which are carefully crafted input images designed to mislead the model into making incorrect predictions [22]. While general, untargeted adversarial attacks only aim to induce incorrect outputs, targeted adversarial attacks present a more insidious threat, manipulating the model’s output to yield an adversary-specified, predetermined response. For instance, a benign image such as a landscape could be subtly altered to elicit harmful text descriptions such as “violence” or “explicit content” from the model. Such manipulation could have severe implications for content moderation systems, potentially leading to the removal of legitimate content or the inappropriate distribution of harmful material.
Figure 1: Comparison of existing targeted adversarial attack strategies and the self-supervised adversarial noise pre-training method AnyAttack with fine-tuning capability.
@@
Both targeted adversarial attacks and jailbreak techniques [38, 1, 2] aim to induce harmful responses from models. However, the key distinction lies in their outputs: targeted adversarial attacks produce adversary-specified, predetermined responses, whereas jailbreak attacks elicit non-predetermined responses. As VLMs become increasingly accessible to the public, facilitating the rapid proliferation of downstream applications, this vulnerability poses a significant threat to the reliability and security of VLMs in real-world scenarios. Therefore, exploring new targeted attack methods tailored to VLMs is crucial to address these vulnerabilities. However, existing targeted attack methods on VLMs present challenges due to the reliance on target labels for supervision, which limits the scalability of the training process. For example, it is impractical to expect a generator trained on ImageNet [19] to produce effective adversarial noise for VLMs.
To overcome this limitation, we propose a novel self-supervised framework, AnyAttack, which leverages the original image itself as supervision, enabling any image to serve as a target for the deployment of targeted adversarial attacks. Our approach involves pre-training a generator on the large-scale LAION-400M dataset [20], enabling the pre-trained noise generator to learn comprehensive noise patterns from diverse image data. Through self-supervised adversarial noise pre-training, we can further fine-tune the pre-trained generator on downstream datasets for adapting downstream vision-language tasks. The large-scale pre-training establishes a foundation model (we release the model weights), thereby enhancing the potential for developing more powerful adversarial attacks. Our framework, to the best of our knowledge, is the first to implement the “pre-training and fine-tuning” paradigm for targeted adversarial attacks at scale. Fig. 1 highlights the distinctions between our method and existing strategies.
@@
To demonstrate the effectiveness of our approach, we conduct extensive experiments on 5 target VLMs (CLIP [18], BLIP [10], BLIP2 [11], InstructBLIP [3], and MiniGPT-4 [37]), across 3 multimodal tasks (image-text retrieval, multimodal classification, and image captioning). We also evaluate our method on commercial VLMs, including Google Gemini, Claude Sonnet, Microsoft Copilot and OpenAI GPT.
@@
  • We propose AnyAttack, a self-supervised framework that utilizes the original image as supervision, allowing any image to serve as a target for deploying targeted adversarial attacks.
  • Our framework is the first to adopt the “pre-training and fine-tuning” paradigm for targeted adversarial attacks, pre-training a noise generator on the large-scale LAION-400M dataset and fine-tuning it for downstream vision-language tasks.
  • We demonstrate the effectiveness of our AnyAttack on five mainstream open-source VLMs across three multimodal tasks. Additionally, we successfully transfer our attack to four commercial VLMs.
Figure 2: Overview of the proposed AnyAttack: a self-supervised framework consisting of pre-training and fine-tuning stages.
@@ 2 Related Work
Targeted Adversarial Attacks.
@@
A number of works have been proposed to improve the effectiveness and transferability of targeted adversarial attacks against vision models. Input augmentation techniques such as image translation [5], cropping [26], mixup [25, 13], and resizing [29], have been employed to increase the diversity of adversarial input, thus improving their transferability across different target models. In addition, adversarial fine-tuning and model enhancement techniques have been explored to increase the attack capabilities of surrogate models [21, 34, 27]. These methods typically involve retraining the surrogate models with a mix of clean and adversarial examples to improve their robustness against future attacks. Furthermore, optimization techniques have evolved to stabilize the update processes during adversarial training. Methods such as adaptive learning rates and gradient clipping have been integrated to ensure more consistent updates and enhance the overall performance of the adversarial attacks [4, 24, 12]. These advancements collectively contribute to the development of more effective and transferable adversarial attacks in the realm of vision models.
Jailbreak Attacks on VLMs.
@@
VLMs have revolutionized DNNs by using large-scale pre-training on diverse image-text datasets. Broadly, VLMs can be categorized into two types: the first offers multimodal functionalities built on large language models (LLMs), complemented by visual models, such as BLIP2 [11], InstructBLIP [3], and MiniGPT-4 [37]. The second type provides a more balanced approach, bridging textual and visual modalities efficiently, as seen in models like CLIP [18], ALIGN [9], and BLIP [10]. Recent advancements in VLMs have spurred parallel research into their vulnerabilities, with jailbreaks and adversarial attacks emerging as distinct threat vectors. Multimodal jailbreaks primarily exploit cross-modal interaction vulnerabilities in VLMs, with the intention of influencing LLMs [38]. These attacks manipulate inputs of text [28], images [2, 8, 17, 16], or both simultaneously [32, 23], aiming to elicit harmful but non-predefined responses. In contrast, image-based adversarial attacks focus on manipulating the image encoder of VLMs, typically leaving the LLM component largely undisturbed. The objective is to induce adversary-specified, predetermined responses through precise visual manipulations. Understanding these differences is crucial for explaining our method.
Adversarial Attacks on VLMs.
@@
Adversarial research on VLMs is relatively limited compared to the extensive studies on vision models, with the majority of existing attacks focusing primarily on untargeted attacks. Co-Attack [33] was among the first to perform white-box untargeted attacks on several VLMs. Following this, more approaches have been proposed to enhance adversarial transferability for black-box untargeted attacks [14, 36, 31, 30]. Cross-Prompt Attack [15] investigates a novel setup for adversarial transferability based on the prompts of LLMs. AttackVLM [35] is the most closely related work, using a combination of text inputs and popular text-to-image models to generate guided images for creating targeted adversarial images. Although their approach shares a similar objective with our work, our method distinguishes itself by being self-supervised and independent of any text-based guidance.
@@ 3 Proposed Attack
In this section, we first present the preliminaries on targeted adversarial attacks and then introduce our proposed method.
3.1 Preliminaries
Threat Model.
@@ 3.1 Preliminaries
We begin by formulating the problem of targeted adversarial attacks. Let $f_{s}$ represent a pre-trained surrogate model, and $\mathcal{D}=\{(x,y)\}$ denote the image dataset, where $x$ is the original image and $y$ is the corresponding label (description). The attacker’s objective is to craft an adversarial example $x^{\prime}=x+\delta$ that misleads the target model $f_{t}$ into predicting a predefined target label $y_{t}$. In the context of VLMs, this objective requires that $x^{\prime}$ aligns with $y_{t}$ as a valid image-text pair. The process of generating targeted adversarial images typically involves finding a perturbation $\delta$ using the surrogate model $f_{s}$.
@@ 3.2 AnyAttack
Our proposed framework, AnyAttack, employs two phases: self-supervised adversarial noise pre-training and self-supervised adversarial noise fine-tuning. Fig. 2 provides the self-supervised framework overview including pre-training and fine-tuning.
@@
For self-supervised adversarial noise pre-training, we train a decoder $F$, to produce adversarial noise $\delta$ on large-scale datasets $\mathcal{D}_{p}$, using frozen encoder $E$ as the surrogate model. Given a batch of images $x$, we extract their embeddings using a frozen image encoder $E$. These normalized embeddings $\mathbf{z}$ are then fed into the decoder $F$, which generates adversarial noise $\delta$ corresponding to the images $x$. To enhance generalization and computational efficiency, we introduce a $K$-augmentation strategy that creates multiple shuffled versions of the original images within each mini-batch. During this process, adversarial noise is added to the shuffled original images (random images) to produce the adversarial images.
For self-supervised adversarial noise fine-tuning, we adapt the pre-trained decoder $F$ to a specific downstream dataset $\mathcal{D}_{f}$. We use an unrelated random image $x_{r}$ from an external dataset $\mathcal{D}_{e}$ as the clean image to synthesize the adversarial image $x_{r}+\delta$.
Self-supervised Adversarial Noise Pre-training
aims to train the generator on large-scale datasets, enabling it to handle a diverse array of input images as potential targets. Unlike existing methods, it does not require target labels or target images as supervision throughout the training process. Our objective can be formulated as follows:
$\min\mathcal{L}(f_{s}(\delta+x_{r}),f_{s}(x)),\quad\text{s.t.}\ x_{r}\neq x,$ (1)
where $x_{r}$ is a random image that is unrelated to $x$, while the adversarial noise $\delta$ is designed to align with the original image $x$ within the surrogate model’s embedding space, and $\mathcal{L}$ denotes the similarity function.
@@
Given a batch of $n$ images $x\in\mathbb{R}^{n\times H\times W\times 3}$ from the large-scale training dataset $\mathcal{D}_{p}$, we employ the CLIP ViT-B/32 image encoder, which is frozen during training, as the encoder $E$, to obtain the normalized embeddings $E(x)=\mathbf{z}\in\mathbb{R}^{n\times d}$ corresponding to the original images $x$, where $d$ represents the embedding dimension (i.e., 512 for CLIP ViT-B/32). Subsequently, we deploy an initialized decoder $F$, which maps the embeddings $\mathbf{z}$ to adversarial noise $D(\mathbf{z})=\delta\in\mathbb{R}^{n\times H\times W\times 3}$ corresponding to the original images $x$. We expect the generated noises $\delta$ to serve as adversarial noise representative of the original images $x$. Our goal is for the generated noises $\delta$, when added to random images $x_{r}$, to be interpreted by the encoder $E$ as the original images $x$, i.e., $E(x_{r}+\delta)=E(x)$.
@@
To increase the number of random images within every batch, we present the $K$-augmentation strategy, which duplicates both adversarial noises $\delta$ and the original images $x$ $K$ times, forming $K$ mini-batches. For each mini-batch, the order of the adversarial noises remain consistent, while the order of the original images is shuffled within the mini-batch, referred to as shuffled images. These shuffled images are then added to the corresponding adversarial noise, resulting in adversarial images $x^{\prime}$. Next, the adversarial images are fed into $F$ to produce adversarial embeddings $\mathbf{z}^{(adv)}$, which are then used for subsequent calculations against the original embeddings $\mathbf{z}$.
Finally, we introduce the adversarial noise pre-training loss $\mathcal{L}_{\text{Pre}}$. It maximizes the cosine similarity between positive sample pairs, defined by the $i$-th elements of adversarial and original embeddings in each mini-batch, while minimizing the similarity between the negative pairs, which consist of all other elements. This setup creates $n$ positive pairs and $n(n-1)$ negative pairs in every mini-batch, with gradients accumulated to update $F$:
$\mathcal{L}_{\text{Pre}}=-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\exp\left(\mathbf{z}_{i}\cdot\mathbf{z}_{i}^{(adv)}/\tau(t)\right)}{\sum_{j=1}^{n}\exp\left(\mathbf{z}_{i}\cdot\mathbf{z}_{j}^{(adv)}/\tau(t)\right)},$ (2)
@@
where $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{(adv)}$ are the $\ell_{2}$-normalized embeddings of the $i$-th sample from original images $x$ and adversarial images $x^{\prime}$. $\tau(t)$ is the temperature at step $t$, enabling the model to dynamically adjust the hardness of negative samples during training. We set a relatively large initial temperature $\tau_{0}$ at the beginning of training and gradually decrease it, reaching the final temperature $\tau_{\text{final}}$ after a certain number of steps $T$:
@@
Self-supervised Adversarial Noise Fine-tuning
refines the pre-trained decoder $F$ on downstream vision-language datasets using task-specific objective functions, facilitating its adaptation to particular domains and multimodal tasks.
Given a batch of $n$ images $x\in\mathbb{R}^{n\times H\times W\times 3}$ from the downstream dataset $\mathcal{D}_{f}$, the encoder $E$ remains frozen and outputs the embeddings $\mathbf{z}$, which are then fed into the decoder $F$ to generate the noise $\delta$. Since the size of $\mathcal{D}_{f}$ is much smaller than that of $\mathcal{D}_{p}$, we randomly select images from an external dataset $\mathcal{D}_{e}$ as random images $x_{r}$, which are then added to the generated noise $\delta$ to create adversarial images. To improve transferability, we incorporate auxiliary models alongside the encoder $E$, forming an ensemble surrogate.
Depending on the downstream tasks, self-supervised adversarial noise fine-tuning employ two different fine-tuning objectives. The first strategy is tailored for the image-text retrieval task, which imposes stricter requirements for distinguishing between similar samples. It demands robust retrieval performance in bi-directions: from $\mathbf{z}^{(adv)}$ to $\mathbf{z}$ and from $\mathbf{z}$ to $\mathbf{z}^{(adv)}$, denoted as $\mathcal{L}_{\text{Bi}}$.
$\mathcal{L}_{\text{Bi}}=\frac{1}{2n}\sum_{i=1}^{n}\Bigg{(}$ $-\log\frac{\exp(\mathbf{z}_{i}\cdot\mathbf{z}_{i}^{(adv)}/\tau)}{\sum_{j=1}^{n}\exp(\mathbf{z}_{i}\cdot\mathbf{z}_{j}^{(adv)}/\tau)}$ (4)
$-\log\frac{\exp(\mathbf{z}_{i}^{(adv)}\cdot\mathbf{z}_{i}/\tau)}{\sum_{j=1}^{n}\exp(\mathbf{z}_{i}^{(adv)}\cdot\mathbf{z}_{j}/\tau)}\Bigg{)}.$